# An exact approach for the multi-depot electric vehicle scheduling problem

### v1

@@ 7 Conclusion
In this work, we have modeled and solved the electric vehicle scheduling problem with multiple depots, multiple charging stations and partial recharge. We have developed a graph representation that is an acyclic network and allows only time-feasible paths. Only two additional aspects need to be ensured for each vehicle schedule: the state-of-charge of the vehicles along a path, ensuring that vehicles cannot run out of energy, and that each vehicle returns to the same depot as it started from.
While our 3-index MILP can be solved with any off-the-shelf solver directly, in order to accommodate multiple depots, constraints of exponential size are incorporated into our 2-index MILP formulation. We compare two types of these constraints and different tailored separation strategies. The best performing strategy relies on infeasible path constraints, separated only at new integer incumbent solutions during the execution of the branch-and-cut algorithm. The 2-index-based branch-and-cut algorithm consistently solves more instances to optimality and in lower computation times than the 3-index model solved by CPLEX, for a low number of depots as well as for six and eight depots. Our approaches have been developed within the collaborative research project ZEMoS (Zero Emission Mobility Salzburg) and serve as decision support for fleet sizing decisions in two pilot regions in the country of Salzburg.
Future work will involve the development of heuristic approaches for more complex problem versions, such as heterogeneous vehicle fleets.
Acknowledgments

### v2

@@ 8 Conclusion
In this work, we have modeled and solved the electric vehicle scheduling problem with multiple depots, multiple charging stations, and partial recharge. We have developed a graph representation that is an acyclic network and allows only time-feasible paths. Only two additional aspects need to be ensured for each vehicle schedule: the SOC of the vehicles along a path, ensuring that vehicles cannot run out of energy, and that each vehicle returns to the same depot as it started from.
While our 3-index MILP can be solved with any off-the-shelf solver directly, in order to accommodate multiple depots, constraints of exponential size are incorporated into our 2-index MILP formulation. We compare two types of these constraints and different tailored separation strategies. The best performing strategy for a lower number of depots relies on infeasible path constraints, separated only at new integer incumbent solutions during the execution of the branch-and-cut algorithm, while for a higher number of depots connectivity cuts, separated at new integer incumbent solutions with setting All works best. The 2-index-based branch-and-cut algorithm consistently solves more instances to optimality and in lower computation times than the 3-index model solved by CPLEX, for a low number of depots as well as for six and eight depots.
The comparison of realistic scenarios reveals important insights in the operational resilience and adaptability of each propulsion technology. Across all scenarios and technologies, we observe that combining multiple lines into a single scheduling problem reduces the total fleet size required. However, the number of recharging/refueling events decrease disproportionately on average, i.e., each individual vehicle charges more than in the separate case.
In general, serving the schedule requires at least as many, and often more, BEBs compared to DBs or FCEBs. BEBs are notably sensitive to environmental constraints, resulting in higher fleet requirements, whereas FCEBs can maintain stable fleet sizes even under cold temperatures. This indicates that FCEBs are more resilient to changes in operational conditions without significant performance degradation.
Interestingly, in the more restrictive cold temperature and battery preservation scenarios, despite requiring larger fleets and significantly more recharging, optimality gaps decrease for both BEBs and FCEBs. This suggests that tighter operational constraints can reduce the solution space and help guide the solver more efficiently toward high-quality solutions.
Taken together, these findings emphasize that integration between lines is a powerful lever to improve operational efficiency, but its implications, particularly regarding vehicle and infrastructure utilization, differ significantly between technologies.
Our approaches have been developed within the collaborative research project ZEMoS (Zero Emission Mobility Salzburg) and serve as decision support for fleet sizing decisions in two pilot regions in the country of Salzburg. Future work will involve the development of heuristic approaches for more complex problem versions, such as heterogeneous vehicle fleets.